OpenStudy (anonymous):

dr miller starts with 3.4 gallons of chlorine in his pool. each day he adds .25 gallons of chlorine and 15% evaporates. how much chlorine will be in his pool at the end of the sixth day

OpenStudy (jack1):

k, so first: set up the equation

OpenStudy (anonymous):

what???

OpenStudy (jack1):

Start amount = 3.4 gal add .25 gal daily 15 % of the total evaporates daily, lets assume after the .25 is added, so multiply by 0.85 daily (0.85 + 0.15 = 1.0, so this will subtract the 15% )

OpenStudy (jack1):

so ur equation = (3.4 + 0.25 ) x 0.85 ... per day, the answer equals the start amount for the next day, (ie replaces the 3.4)...

OpenStudy (jack1):

\(\large x = (((((((3.4 + 0.25) \times 0.85) +0.25) \times 0.85) +0.25) \times 0.85) +0.25) \times 0.85) +0.25) \times 0.85) +0.25)\)

OpenStudy (jack1):

i got day 6 as 2.87698 gal, hbu?